# LLäMmlein [Uncaptioned image]: Transparent, Compact and Competitive German-Only Language Models from Scratch

### v2

@@ 4.3.2 LLäMmlein 1B
Table 4: Performance of our models and instruction tuned variants on the lm-evaluation-harness-de, with TruthfulQA (mc2), ARC-Challenge (acc_norm), HellaSwag (acc_norm), MMLU (acc). Short version of Table 8.
@@
For smaller models, Llama 3.2 (1B) achieved the best TruthfulQA score. However, in completion tasks like ARC-Challenge and HellaSwag, LLäMmlein 1B Instruct models consistently outperformed both the base LLäMmlein model and Llama 3.2 1B, indicating that chat-tuning enhances performance on structured completion tasks. Conversely, Llama 3.2 1B Instruct excelled in the broader knowledge-focused MMLU benchmark. Interestingly, chat-tuning improved LLäMmlein scores across all tasks, a trend not observed for the Llama 3.2 model.
@@ 4.4 Checkpoint Averaging
Checkpoint averaging did not improve - or even change - downstream task performance on SuperGLEBer for either the 120M or 1B model (see Figure 9 for the 1B model). This was unexpected, but we hypothesize the checkpoints being too far apart, as Vaswani et al. (2017) averaged checkpoints written every 10 minutes near the end of training, while our checkpoints are about 6-8 hours apart.
[Uncaptioned image] Interesting: Checkpoint averaging ineffective, possibly checkpoints are too far apart
@@ 7 Limitations
While the LLäMmlein models represent a significant contribution to German NLP research, several limitations remain: 1. Lack of Granular Training Sample TrackingUnlike approaches such as Pythia, when exactly the inclusion of individual samples in the training process is not explicitly traceable, which could enhance reproducibility and error analysis. 2. Limited Capabilities on some domainsDue to the scarcity of high-quality German resources for e.g. coding, we found the models perform poorly on such tasks. 3. Monolingual FocusWhile being considered a strength in the context of this setup, LLäMmlein lacks the ability to leverage multilingual contexts or perform cross-lingual tasks, which could limit usability in certain scenarios. 4. Evaluation ScopeWhile evaluated extensively on the SuperGLEBer benchmark and lm-evaluation-harness, other domains such as literature, spoken language, or dialects were not tested, leaving gaps in the understanding of model capabilities. 5. Long-Context HandlingThe models were trained with a maximum sequence length of 2048 tokens, which limits their applicability to tasks requiring extended contexts, such as processing long documents or legal texts.
@@ Acknowledgments
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b185cb. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Furthermore, we gratefully acknowledge the HPC resources provided by the JuliaV2 cluster at the Universität Würzburg (JMU), which was funded as DFG project as “Forschungsgroßgerät nach Art 91b GG” under INST 93/1145-1 FUGG. The project staff is partially funded by the DFG - 529659926. The data science chair is part of the CAIDAS, the Center for Artificial Intelligence and Data Science, and is supported by the Bavarian High-Tech Agenda, which made this research possible.
@@ Appendix D Training
Architectural and training details for both LLäMmlein models can be found in Table 6.
Table 6: Architectural and training details of LLäMmlein models
@@
In addition, we provide the loss curves for both models in Figures 3 and 4.
@@
Figure 4: Loss curve of LLäMmlein 1B model. Each color indicates a run, resumed after a training interruption. The visible jumps in the loss curve correspond to different chunks of training data, each sampled from a distinct part of the dataset.
For LLäMmlein-120M overall, 10 restarts were necessary: Due to cluster settings, training was resumed at least every two days, and additionally, training had to be restarted a few times to address GPU and NCCL errors. Before starting the training run for LLäMmlein-1B, we preliminary attempted to estimate the runtime for replicating the original TinyLlama training settings on our hardware as a sanity check. Based on our extrapolations, the process would have taken over 200 days using 16 A100 GPUs, compared to the 90 days reported for TinyLlama on the same hardware. After suspecting infiniband/sharding configuration issues, we adapted the Fully Sharded Data Parallel (FSDP) strategy to a hybrid sharding approach. This reduced the extrapolated runtime to approximately 100 days, which we deemed satisfactory. Next, we scaled our training to the final 64 GPUs, bringing the extrapolated runtime down to 36 days. Early in the training run, we identified further inefficiencies related to improper use of the available InfiniBand. We halted the training, corrected the configuration, switched back to full sharding, and implemented dataset pre-caching in RAM on each node. After these optimizations, we restarted the training, achieving a final overall runtime of 32 days (already including a few slower initial days before the final configuration change (green in Figure 4)), with a total of two restarts, illustrated in different colors in Figure 4. We will publish the code including all mentioned fixes/adaptations upon publication.
@@ Appendix E Evaluation on SuperGLEBer
Figure 7 illustrates comparisons (incl. t-tests) of our model against existing models of the same size on the SuperGLEBer benchmark. LLäMmlein clearly outperforms german-gpt2, confirming its superiority among German decoder models of similar size. When comparing LLäMmlein with the two BERT models - gbert-base and bert-base-german-cased - no statistically significant differences were found, which makes our 120M model the first to match the average performance of a similarly sized encoder on the SuperGLEBer benchmark.
@@
Table 8: Performance comparison of LLäMmlein and its Instruction tuned variants as well as similar sized and various larger models on the lm-evaluation-harness-de including the: TruthfulQA (mc1|mc2), ARC-Challenge (acc|acc_norm), HellaSwag (acc|acc_norm) and MMLU (acc|acc_norm). This is the full table of Table 4, including all metrics and all trained chat adapters.
@@
Table 10: Example outputs in German, Bavarian, and Swiss dialects for the base model and two adapter variants, when being asked about the “regional captial”, showcasing linguistic adaptability across regional varieties. The outputs demonstrate varying levels of accuracy in adapting to the target dialects. While some results capture the linguistic characteristics effectively, others fall short, highlighting the challenges of fine-tuning models for nuanced regional variations. The tendency of the bavarian adapter to adhere to the Wikipedia is obvious as it starts with a short descriptive sentence every time.

### v3

@@ 4.3.2 LLäMmlein 1B
Table 4: Performance of our (instruction tuned) models on the lm-evaluation-harness-de, with TruthfulQA (mc2), ARC-Challenge (acc_norm), HellaSwag (acc_norm), MMLU (acc). Short version of Table 8.
@@
For smaller models, Llama 3.2 (1B) achieved the best TruthfulQA score. However, in completion tasks like ARC-Challenge and HellaSwag, LLäMmlein 1B Instruct models consistently outperformed both the base LLäMmlein model and Llama 3.2 1B, indicating that instruct-tuning enhances performance on structured completion tasks. Conversely, Llama 3.2 1B Instruct excelled in the broader knowledge-focused MMLU benchmark. Interestingly, instruct-tuning improved LLäMmlein scores across all tasks, a trend not observed for the Llama 3.2 model.
@@ 4.4 Checkpoint Averaging
Checkpoint averaging did not improve – or even change – downstream task performance on SuperGLEBer for either the 120M or 1B model (see Figure 9 for the 1B model). This was unexpected, but we hypothesize the checkpoints being too far apart, as Vaswani et al. (2017) averaged checkpoints written every 10 minutes near the end of training, while our checkpoints are about 6-8 hours apart.
[Uncaptioned image] Interesting: Checkpoint averaging ineffective, possibly checkpoints are too far apart.
@@ 7 Limitations
While the LLäMmlein models represent a significant contribution to German NLP research, several limitations remain: 1. Limited Capabilities on some domainsDue to the scarcity of high-quality German resources for e.g. coding, we found the models perform poorly on such tasks. 2. Monolingual FocusWhile being considered a strength in the context of this setup, LLäMmlein lacks the ability to leverage multilingual contexts or perform cross-lingual tasks, which could limit usability in certain scenarios. 3. Evaluation ScopeWhile evaluated extensively on the SuperGLEBer benchmark and lm-evaluation-harness, other domains such as literature, spoken language, or dialects were not tested, leaving gaps in the understanding of model capabilities. 4. Long-Context HandlingThe models were trained with a maximum sequence length of 2048 tokens, which limits their applicability to tasks requiring extended contexts, such as processing long documents or legal texts.
@@ Acknowledgments
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b185cb. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Furthermore, we gratefully acknowledge the HPC resources provided by the JuliaV2 cluster at the Universität Würzburg (JMU), which was funded as DFG project as “Forschungsgroßgerät nach Art 91b GG” under INST 93/1145-1 FUGG. The project staff is partially funded by the DFG – 529659926. The data science chair is part of the CAIDAS, the Center for Artificial Intelligence and Data Science, and is supported by the Bavarian High-Tech Agenda, which made this research possible. We would also like to thank Anton Vlasjuk for his valuable contributions and insights, particularly for his efforts in integrating Flash Attention.
@@ Appendix D Training
Table 6: Architectural and training details of LLäMmlein models
Architectural and training details for both LLäMmlein models can be found in Table 6. In addition, we provide the loss curves for both models in Figures 3 and 4.
@@
Figure 4: Loss curve of LLäMmlein 1B model. Each color indicates a run, resumed after an interruption. The visible jumps correspond to different chunks of training data, each sampled from a distinct part of the dataset.
For LLäMmlein-120M overall, ten restarts were necessary: Due to cluster settings, training was resumed at least every two days, and additionally, training had to be restarted a few times to address GPU and NCCL errors. Before starting the training run for LLäMmlein-1B, we preliminary attempted to estimate the runtime for replicating the original TinyLlama training settings on our hardware as a sanity check. Based on our extrapolations, the process would have taken over 200 days using 16 A100 GPUs, compared to the 90 days reported for TinyLlama on the same hardware. After suspecting sharding configuration issues, we adapted the Fully Sharded Data Parallel (FSDP) strategy to a hybrid sharding approach. This reduced the extrapolated runtime to approximately 100 days, which we deemed satisfactory. Next, we scaled our training to the final 64 GPUs, bringing the extrapolated runtime down to 36 days. Early in the training run, we identified further inefficiencies related to improper use of the available InfiniBand. We halted the training, corrected the configuration, switched back to full sharding, and implemented dataset pre-caching in RAM on each node. After these optimizations, we restarted the training, achieving a final overall runtime of 32 days (already including a few slower initial days before the final configuration change (green in Figure 4)), with a total of two restarts, illustrated in different colors in Figure 4. We will publish the code including all mentioned fixes/adaptations upon publication.
@@ Appendix E Evaluation on SuperGLEBer
Figure 7 illustrates comparisons (incl. t-tests) of our model against existing models of the same size on the SuperGLEBer benchmark. LLäMmlein clearly outperforms german-gpt2, confirming its superiority among German decoder models of similar size. When comparing LLäMmlein with the two BERT models – gbert-base and bert-base-german-cased – no statistically significant differences were found, which makes our 120M model the first to match the average performance of a similarly sized encoder on the SuperGLEBer benchmark.
@@
Table 8: Performance comparison of (instruction tuned) LLäMmlein variants as well as similar sized and various larger models on the lm-evaluation-harness-de including the: TruthfulQA (mc1|mc2), ARC-Challenge (acc|acc_norm), HellaSwag (acc|acc_norm) and MMLU (acc|acc_norm). This is the full table of Table 4, including all metrics and all trained instruct adapters.
@@
Table 10: Example outputs in German, Bavarian, and Swiss dialects for the base model and two adapter variants, when being asked about the “regional captial”, showcasing linguistic adaptability across regional varieties. The outputs demonstrate varying levels of accuracy in adapting to the target dialects. While some results capture the linguistic characteristics effectively, others fall short, highlighting the challenges of fine-tuning models for nuanced regional variations. The tendency of the Bavarian adapter to adhere to the Wikipedia style is pronounced as it starts with a short descriptive sentence every time.